# UniDoc-Bench: A Unified Benchmark for Document-Centric Multimodal RAG

### v1

@@ 5 Conclusion
Reproducibility Statement
We ensure the reproducibility of our work. All source datasets we employed are publicly available (PDFA). We are making our code available in the supplementary materials to enable replication of our findings.
Acknowledgements
We would like to sincerely thank Nesrine Yakoubi, Caitlyn Cline, Wyatt Miller, Michael Thuo, John Soledad, and Fabriana Louisa Pita for their invaluable efforts in annotating the datasets. Their careful work and dedication were essential to the success of this research.
@@ B.5.1 Text-only
Prompt P.1: Text-only RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: The 2-hop synthesized question must be a single, self-contained question and must not use ”and” to connect multiple questions. The answer of the synthesized question will only be found in the contexts. The answer of the synthesized question cannot be found in the images. The synthesized question must require all the chunks in the contexts to be answered. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. You must also provide an explanation why the answer can only be found in the provided contexts. Question Template: Use the following template to generate the QA:
@@ B.5.2 Image-only
Prompt P.2: Image-only RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: 1. The synthesized question must be a single, self-contained question and must not use “and” to connect multiple questions. 2. The answer of the synthesized question will only be found in the image and cannot be found in any sentences in the chunks of the provided contexts. 3. The synthesized question must require chunks/contexts to locate the image and cannot mention the image directly. 4. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. 5. Do not ask “what XYZ in the graph/image/figure”; the question must be general enough to be asked in a large corpus. 6. If you cannot synthesize a question which can only be answered in the image based on the above requirements, do not synthesize anything. 7. Provide an explanation why the answer can only be found in the image and cannot be found in the provided chunks/contexts. 8. Avoid phrasing like “what is shown in the image,” e.g., ”what color/logo/name in the image.” 9. Emphasize reasoning, aggregation, temporal comparison, or retrieval from source data. Imagine the question being asked without the image still making partial sense. Question Template: Use the following template to generate the QA:
@@ B.5.3 Text-plus-Image
Prompt P.3: Text-plus-image RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: 1. The 2-hop synthesized question must require both the provided contexts and images to answer. 2. The concise answer of the synthesized question will directly require information in the image to answer. 3. The concise answer of the synthesized question will also require information in the natural language contexts to answer. 4. The synthesized question must require contexts to locate the image and cannot mention the image directly. 5. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. 6. Provide an explanation indicating which part of the image is used to answer and which sentence in the contexts is used to answer the question. 7. Do not ask “what XYZ in the graph”; the question must be general enough to be asked in a large corpus. 8. If you cannot synthesize a question based on these requirements or directly use the information in the images, do not synthesize anything. 9. If the image can only be used for visualization or illustration, do not synthesize anything. If you cannot use all the chunks in the answer, do not synthesize the question. 10. The synthesized question must be a single, self-contained question and must not use “and” to connect multiple questions. Question Template: Use the following template to generate the QA:
@@ B.6 Rewriting prompts
Example 1: Original: “What is the revenue growth shown in Figure 3 in 2024’s report?”
@@
Example 2: Original: “What is the median differential rate between hurdle rates and costs of capital for cyclical and non-cyclical firms?”
@@ B.7 Answer Rewriting Prompts
Input Data: Question: “{{rewritten_question_obscured}}” Contexts: “{{contexts}}” Old Correct Answer: “{{answer}}” Images: The image is as follows:
@@ C.2.2 Label Definitions
Note: The original instructions included a rule stating, ”If a question or response is not factual, it should be labeled as ‘Incomplete’.” However, the provided examples use the ”Not Factual” label, which was the standard followed during annotation.
@@ C.5.2 Label Definitions
The question is self-contained. It is clearly phrased, makes sense on its own, and provides enough specific detail (e.g., names, topics, concepts) to be answerable. It does not rely on vague document references. For example, ”What are the key benefits of solar energy mentioned in the 2022 Department of Energy report?” is self-contained.
@@
The question depends on external or implicit context to be meaningful. It may contain vague deictic references (e.g., ”in the image above,” ”according to this chart,” ”what does this mean?”) without clarifying what the reference points to. For example, ”What is the logo in the image?” is not self-contained as it requires seeing a specific, un-referenced image.
@@
Listing 6: Example of a question that is not self-contained due to a vague reference (”the image”).
@@
Listing 7: Example of a question that is self-contained because it provides sufficient context (”personalization strategies,” ”telecommunications company”).
@@ C.6.2 Label Definitions
The question represents a reasonable and natural query a human would make. It seeks meaningful information such as facts, summaries, comparisons, or explanations, and is phrased in a way that reflects a real information need. For example: ”What were the company’s main revenue streams in the last fiscal year?”
@@ Appendix E Additional Experiments
E.1 Retrieval Performance
We break down retrieval performance by question and answer types, as reported in Table 7. We find that question type has minimal impact on retrieval recall, whereas answer type plays a significant role. For text-only retrieval, performance is substantially higher on questions requiring text to answer, but markedly lower on image-required questions. Conversely, for image-only retrieval, questions requiring image-based answers are retrieved more effectively than those requiring text, highlighting the modality-specific strengths of each embedding approach. Combining both embeddings (T++I) effectively leverages the advantages of each modality, resulting in higher overall recall. For multimodal embeddings, image-required questions tend to be retrieved more easily than text-required questions, suggesting that current multimodal embeddings function more like image retrieval in practice.
Table 7: Retrieval performance (Precision@10 / Recall@10) of four RAG systems on 1600 QA pairs, averaged across eight domains and broken down by question and answer types.
Type Text (OpenAI) IMG ( colqwen) MM (GME) T++I
Prec. Recall Prec. Recall Prec. Recall Prec. Recall
Factual Retrieval 0.319 0.759 0.237 0.839 0.304 0.876 0.416 0.862
Comparison 0.440 0.839 0.276 0.854 0.368 0.901 0.503 0.891
Summary 0.497 0.856 0.329 0.830 0.400 0.907 0.563 0.883
Logical 0.496 0.801 0.306 0.789 0.381 0.832 0.537 0.829
Text-only 0.511 0.821 0.324 0.774 0.390 0.836 0.558 0.820
Img-only 0.152 0.751 0.174 0.922 0.258 0.900 0.273 0.916
Text + Img 0.489 0.850 0.306 0.833 0.392 0.907 0.555 0.880
Table-required 0.431 0.773 0.270 0.798 0.339 0.872 0.493 0.851
E.2 MM-embedding RAG Comparison
We compare RAG performance using two multimodal embeddings: voyage-multimodal-3 and gme-Qwen2-VL-7B-Instruct, with results reported in Table 8 and Table 9. While voyage-multimodal-3 achieves slightly lower recall but higher precision in retrieval compared to gme-Qwen2-VL-7B-Instruct, it delivers better overall performance when integrated into MM-RAG.
Table 8: Retrieval performance (Precision@10 / Recall@10) and end-to-end performance (Recall using retrieved-top-10 and retrieved-top-20 candidates) of two MM-RAG systems on 200 QA pairs across eight domains, with average recall reported across all domains.
Domain MM (Voyage) MM (GME)
Retrieval End-to-end Retrieval End-to-end
Prec. Recall top-$10$ top-2020 Prec. Recall top-$10$ top-2020
Commerce 0.518 0.892 0.629 0.653 0.354 0.895 0.617 0.611
Construction 0.406 0.733 0.603 0.609 0.336 0.881 0.601 0.616
CRM 0.418 0.748 0.634 0.653 0.343 0.884 0.623 0.637
Education 0.419 0.784 0.652 0.658 0.366 0.912 0.640 0.668
Energy 0.418 0.783 0.659 0.680 0.331 0.847 0.669 0.666
Finance 0.426 0.726 0.622 0.644 0.370 0.898 0.627 0.636
Healthcare 0.388 0.766 0.638 0.668 0.376 0.857 0.642 0.664
Legal 0.431 0.764 0.631 0.669 0.327 0.876 0.609 0.629
Avg. 0.416 0.777 0.633 0.654 0.350 0.881 0.628 0.641
Table 9: Precision and recall of two MM-RAG systems using the top $10$ retrieved chunks retrieved by their retrievers, evaluated across different question and answer types on $1,600$ QA pairs spanning eight domains, with average recall reported across all domains.
Type MM (Voyage) MM (GME)
Prec. Recall Prec. Recall
Factual Retrieval 0.606 0.595 0.691 0.580
Comparison 0.656 0.604 0.730 0.608
Summary 0.694 0.738 0.802 0.655
Logical Reasoning 0.699 0.727 0.837 0.679
Text-only 0.871 0.824 0.868 0.759
Img-only 0.414 0.348 0.436 0.312
Text+Img 0.786 0.656 0.810 0.636
Table-required 0.832 0.736 0.867 0.750
E.3 Cost Comparison
We also calculate the average inference cost and latency of different RAG systems. The image-only system (IMG) is the most efficient, while multimodal systems (MM) are the slowest, reflecting the trade-off between complexity and capability. The text-only system consumes the most tokens and is therefore the most expensive. The T+I fusion RAG retrieves from text chunks first, then images, which increases latency. These results suggest that modern MM-RAG systems can offer both improved performance and lower cost compared to text-only RAG.
@@ F.1 Content-rich images increase difficulty
We analyzed all images in the documents of the easiest domain (commerce manufacturing and legal) and the most difficult domains (finance and construction). Using gemini-2.5-pro, we classified images as either content-rich (providing information not present in the text) or illustrative. In finance and construction, 62.8% and 69.3% of images, respectively, were content-rich, compared to 40.0% in commerce manufacturing and 49.5% in legal. This suggests that domains with a higher proportion of content-rich images present a greater challenge for RAG, as these images require effective multimodal understanding beyond text.

### v3

@@ 5 Conclusion
Limitations
UniDoc-Bench relies on LLM-synthesized, template-based queries which, despite human verification, may lack the linguistic diversity and conversational dependency (e.g., multi-turn follow-ups) characteristic of organic user interactions. The evaluation protocol relies on assumptions, such as treating page-level retrieval matches as correct—potentially inflating recall for dense documents—and explicitly excluding uncaptioned figures under the assumption they are non-informative. Furthermore, the benchmark is currently limited to English-centric documents across eight specific domains and employs LLM-based judges for end-to-end metrics, suggesting that findings may not generalize to low-resource languages or remain robust against inherent judge model biases.
@@ B.5.1 Text-only
Prompt P.1: Text-only RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: The 2-hop synthesized question must be a single, self-contained question and must not use "and" to connect multiple questions. The answer of the synthesized question will only be found in the contexts. The answer of the synthesized question cannot be found in the images. The synthesized question must require all the chunks in the contexts to be answered. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. You must also provide an explanation why the answer can only be found in the provided contexts. Question Template: Use the following template to generate the QA:
@@ B.5.2 Image-only
Prompt P.2: Image-only RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: 1. The synthesized question must be a single, self-contained question and must not use “and" to connect multiple questions. 2. The answer of the synthesized question will only be found in the image and cannot be found in any sentences in the chunks of the provided contexts. 3. The synthesized question must require chunks/contexts to locate the image and cannot mention the image directly. 4. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. 5. Do not ask “what XYZ in the graph/image/figure"; the question must be general enough to be asked in a large corpus. 6. If you cannot synthesize a question which can only be answered in the image based on the above requirements, do not synthesize anything. 7. Provide an explanation why the answer can only be found in the image and cannot be found in the provided chunks/contexts. 8. Avoid phrasing like “what is shown in the image,” e.g., "what color/logo/name in the image." 9. Emphasize reasoning, aggregation, temporal comparison, or retrieval from source data. Imagine the question being asked without the image still making partial sense. Question Template: Use the following template to generate the QA:
@@ B.5.3 Text-plus-Image
Prompt P.3: Text-plus-image RAG Question Generation Prompt: You are an assistant specialized in creating Multimodal RAG tasks. The task is the following: Given some natural language contexts and images inside these contexts, you will generate questions that can be asked by a user to retrieve information from a large documentary corpus. Requirements: 1. The 2-hop synthesized question must require both the provided contexts and images to answer. 2. The concise answer of the synthesized question will directly require information in the image to answer. 3. The concise answer of the synthesized question will also require information in the natural language contexts to answer. 4. The synthesized question must require contexts to locate the image and cannot mention the image directly. 5. The synthesized question must be specific enough to locate the contexts in a large documentary corpus. 6. Provide an explanation indicating which part of the image is used to answer and which sentence in the contexts is used to answer the question. 7. Do not ask “what XYZ in the graph"; the question must be general enough to be asked in a large corpus. 8. If you cannot synthesize a question based on these requirements or directly use the information in the images, do not synthesize anything. 9. If the image can only be used for visualization or illustration, do not synthesize anything. If you cannot use all the chunks in the answer, do not synthesize the question. 10. The synthesized question must be a single, self-contained question and must not use “and" to connect multiple questions. Question Template: Use the following template to generate the QA:
@@ B.6 Rewriting prompts
Example 1: Original: “What is the revenue growth shown in Figure 3 in 2024’s report?"
@@
Example 2: Original: “What is the median differential rate between hurdle rates and costs of capital for cyclical and non-cyclical firms?"
@@ B.7 Answer Rewriting Prompts
Input Data: Question: “{{rewritten_question_obscured}}" Contexts: “{{contexts}}" Old Correct Answer: “{{answer}}" Images: The image is as follows:
@@ C.2.2 Label Definitions
Note: The original instructions included a rule stating, "If a question or response is not factual, it should be labeled as ‘Incomplete’." However, the provided examples use the "Not Factual" label, which was the standard followed during annotation.
@@ C.5.2 Label Definitions
The question is self-contained. It is clearly phrased, makes sense on its own, and provides enough specific detail (e.g., names, topics, concepts) to be answerable. It does not rely on vague document references. For example, "What are the key benefits of solar energy mentioned in the 2022 Department of Energy report?" is self-contained.
@@
The question depends on external or implicit context to be meaningful. It may contain vague deictic references (e.g., "in the image above," "according to this chart," "what does this mean?") without clarifying what the reference points to. For example, "What is the logo in the image?" is not self-contained as it requires seeing a specific, un-referenced image.
@@
Listing 6: Example of a question that is not self-contained due to a vague reference ("the image").
@@
Listing 7: Example of a question that is self-contained because it provides sufficient context ("personalization strategies," "telecommunications company").
@@ C.6.2 Label Definitions
The question represents a reasonable and natural query a human would make. It seeks meaningful information such as facts, summaries, comparisons, or explanations, and is phrased in a way that reflects a real information need. For example: "What were the company’s main revenue streams in the last fiscal year?"
@@ Appendix E Additional Experiments
E.1 Cost Comparison
@@
MM-RAG systems can achieve both better end-to-end performance and lower cost than text-only RAG. Text-only RAG is the most expensive due to high token consumption, while image-only RAG has the lowest cost and latency. Multimodal RAG offers lower cost than text-only RAG with comparable latency.
We report the average inference cost and latency of different RAG systems in Table 6. Image-only systems (IMG) are the most efficient, whereas multimodal systems (MM) are the slowest, reflecting the trade-off between model complexity and capability. The T+I fusion RAG incurs additional latency because it retrieves text chunks before images. Overall, these results show that modern MM-RAG systems can provide improved performance at lower cost than text-only RAG.
@@ F.1 Content-rich images increase difficulty
We analyze images from the easiest domains (commerce manufacturing and legal) and the most challenging domains (finance and construction). Using gemini-2.5-pro, we classify images as content-rich (containing information not present in the text) or illustrative. Content-rich images are substantially more common in finance (62.8%) and construction (69.3%) than in commerce manufacturing (40.0%) and legal (49.5%). This suggests that domains with a higher proportion of content-rich images pose greater challenges for RAG, as they require effective multimodal understanding beyond text, consistent with the results in Table 4.